OpenStudy (anonymous):

Find an equation for the nth term of the arithmetic sequence. a19 = -58, a21 = -164

OpenStudy (anonymous):

an = 896 - 53(n - 2) an = 896 - 53(n - 1) an = 896 + 53(n + 1) an = 896 - 53(n + 1)

OpenStudy (anonymous):

@cwrw238 @lgbasallote @Zekarias

OpenStudy (anonymous):

\[a_{19}=a_{1}+18d ....and..a_{21}=a_{1}+20d\]Now find \[a_{1}..and..d\]You will get the answer\[a_{n}=896-53(n-1)\]